### v4

@@ 3.3 Diversity, Debiasing and Problem Setting
Diversity of Questions. The limited diversity in generating questions using pre-designed templates may not significantly impact the significance of datasets for several reasons. First, in the context of sports, there is a high level of restriction imposed by rules and specific content, resulting in a more focused attention from the audience on particular actions and events. This inherent structure in sports content narrows down the range of potential questions, making template-based question generation a feasible approach. Secondly, the practice of template-based question generation is widely accepted and employed in the construction of VideoQA datasets [15, 41, 43]. The adoption of this approach in reputable datasets attests to its effectiveness and acceptance. Therefore, while the questions are template-based, their relevance and significance in the context of sports are preserved, ensuring the robustness of the dataset for analysis and evaluation.
@@
Problem Setting. Following the approach in [15, 41], we define an open-ended task based on the generated QA pairs. It is important to note that the open-ended setting, along with the multiple-choice setting, is widely used in the field of VideoQA. We have chosen to commit to the open-ended task instead of the multiple-choice one for two reasons: 1) The open-ended task is more challenging as it requires models to choose from a large answer set rather than selecting from several given options. 2) Limiting the answer choices to only several options would significantly decrease the diversity of the dataset. Specifically, all answers in the QA pairs form an answer pool, which is treated as 191 classes (after discarding classes with fewer than 30 samples).
@@ 4 Auto-Focus Transformer for Sports Video Question Answering
A key characteristic of the sports VideoQA task is the need for reasoning over temporal dependencies of various scales for multiple objects. For instance, a question asking about the number of actions requires the model to capture global dependencies in the video, whereas a question querying the effect of an action relies on short-term temporal information. The scale of temporal dependency required varies based on the question. However, current Transformer-based or GNN-based VideoQA methods tend to focus on global dependencies, irrespective of the required scale of temporal information. To address this limitation, we propose the Auto-Focus Transformer (AFT), a new Transformer encoder featuring a novel multi-head Auto-Focus Attention (AFA) mechanism, designed specifically for sports VideoQA.
@@
Specifically, we start by extracting the appearance feature and motion feature of the frame using pretrained ResNet [13] and I3D [4] (refer to Section 5.1 for more details). Subsequently, we concatenate these features and input them into a linear layer for dimension reduction and feature fusion. Consequently, for each video, we obtain a sequence of frame representations ($\bm{R}$). Regarding the question, we utilize word embedding and a bi-directional RNN to obtain the text representations, encompassing both the global one ($\bm{w}$) and the local ones ($\bm{W}$).
We then capture the dependencies among frames using AFT where the AFA mechanism is computed as follows (the multi-head AFA is defined likewise): given the key set $\left\{\bm{k}_{i}\right\}$, the value set $\left\{\bm{v}_{i}\right\}$, and a query $\bm{q}_{j}$ ($\bm{q}_{j},\bm{k}_{i},\bm{v}_{i}\in\mathbb{R}^{d}$ are mapped from $\bm{R}$ in AFA),
@@
where $\mathcal{F}$ is a set of pre-defined focal lengths. $\mathcal{D}_{j}^{f}=\{i||i-j|\leq f\}$ is the index set of the neighbours of $\bm{q}_{j}$ within focal length $f$. $\alpha_{f}\in(0,1)$ is the dynamic weight of focus $f$, satisfying $\sum_{f}\alpha_{f}=1$ and depended on the question. As shown in Fig. 5, $\bm{\alpha}=\{\alpha_{f}\}_{f=1}^{\mathcal{|F|}}$ are obtained by applying linear projection and softmax to the global representation of question. Compared to traditional attention that models global dependencies in any case, our AFA captures only the temporal dependencies of the scale required by the question. This design allows the model to focus solely on crucial information and discard irrelevant parts for better prediction. We denote the output of AFT as $\bm{M}$.
For the video-text integration, we draw inspiration from [24] and [14], employing a graph convolutional network (GCN) and block fusion [2]. Refer to [24] for detailed information. Ultimately, we obtain a fused feature $\bm{s}$ and predict the answer using a linear projection and softmax.
@@ 5.1 Experimental Settings
Pre-processing and Representation. The videos in our dataset are down-sampled to 5 FPS to reduce temporal redundancy. We extract two types of visual features for experiments: appearance feature and motion feature. For each feature type, we obtain both global and local features. Specifically, we detect 10 objects in every two frames of each video using Faster-RCNN [30] (with ResNet-101 [13] as the backbone and a score threshold set to 0.6), pretrained on COCO [21]. These detected objects serve as local appearance features in our work. Additionally, we use the global average pooling of Res4 activation in Faster-RCNN as the global appearance feature. To extract motion features, we employ I3D [4] pretrained on Kinetics [4]. Specifically, for each frame, we combine it with its adjacent 7 frames, creating a clip that we then feed to I3D to extract a 3D feature map. The global motion feature of a frame is obtained by applying global average pooling to the 3D feature. The local motion feature of an object in a frame is obtained by applying RoIAlign [12] (output size $1\times 1$) to the 3D feature with its corresponding bounding box. Note that our method uses only the global feature. For language, we explore two types of word embeddings: GloVe [28] and pre-trained BERT [6].
Implementation Details. We set the attention focus to =$\mathcal{F}=\{3,9,80\}$ to cover short-term, mid-term, and long-term dependencies, respectively. The hidden dimension $d$ is set to 512. The loss function is cross-entropy. The models are trained for 50 epochs with a learning rate of $1\times 10^{-4}$ and a batch size of 16, using the Adam optimizer [16].
Evaluation Metrics. For evaluation metrics, we employ accuracy to showcase the models’ performance on our dataset. Additionally, recognizing the long-tail distribution of our dataset, we evaluate methods using F1-score. F1-score is particularly important for sports VideoQA as it explicitly reflects performance averaged over imbalanced classes.
Table 4: Baseline results on Sports-QA, including random choosing, semantic-aware (S-A) random choosing, LSTM-based BlindQA, and Transformer-based BlindQA. The best and the second-best results (%) are bold and underlined.
@@
Random S-A Random LSTM Transformer LSTM Transformer
Accuracy 0.5 24.6 44 43.7 43.9 43.9
F1-score 0.3 5.4 13.5 14.9 13.8 15.8
@@
CoMem [10] 72.8 63.1 69.8 70.0 50.3 77.0 32.9 50.0 60.9 57.1 23.1
HME [8] 64.8 63.3 68.5 71.6 49.3 76.5 31.2 50.4 54.3 56.0 23.6
HGA [14] 71.3 65.4 72.2 72.2 51.3 78.1 34.4 52.0 59.2 58.3 24.5
HQGA [40] 72.8 64.6 70.5 71.6 51.0 77.6 33.8 54.0 58.2 57.9 25.0
IGV [24] 68.8 66.0 70.6 72.4 52.1 78.2 34.7 52.1 59.2 58.5 24.8
MASN [31] 71.2 64.2 70.1 66.9 50.5 76.4 33.6 50.2 58.9 57.0 23.4
@@
CoMem [10] 74.1 64.0 68.7 70.7 49.3 77.2 32.1 48.4 55.6 56.6 26.4
HME [8] 64.7 63.4 66.2 70.6 49.3 76.6 30.7 46.6 53.6 55.6 22.9
HGA [14] 71.6 66.5 69.6 70.7 51.3 78.0 33.8 54.3 56.6 58.1 25.1
HQGA [40] 74.1 63.5 67.2 70.8 50.3 76.6 33.6 49.1 57.9 57.0 25.0
IGV [24] 72.0 63.9 70.4 72.4 51.7 78.0 34.2 51.9 63.2 58.2 23.8
MASN [31] 73.0 63.0 69.7 66.6 50.1 76.4 32.5 50.6 61.2 56.6 24.1
@@
CoMem [10] CVPR’18 Mem
HME [8] CVPR’19 Mem
HGA [14] AAAI’20 GNN
IGV [24] CVPR’22 IG
HQGA [40] AAAI’22 HL
MASN [31] ACL’21 GNN
@@ 5.2 Baseline Study
BlindQA: Models are constructed with a question encoder and an MLP-based answer decoder, without exploiting visual information. Two types of word embeddings (Glove and pre-trained BERT) and two question encoders (LSTM and Transformer [36]) are studied.
The results of the baselines are shown in Table 4. Random choosing exhibits poor accuracy and F1-score, but considering constraints from the questions significantly improves performance. BlindQA achieves considerable results in both accuracy and F1-score, indicating noticeable semantic correlations between questions and answers. By maximizing the likelihood of the answer conditioned on a specific question type, the model fits the answer distribution given different questions in the training set.
@@ 5.3 Benchmark on Sports-QA
We benchmark several VideoQA methods on our Sports-QA dataset, comparing them to CoMem [10], HME [8], HGA [14], MASN [31], HQGA [40], and IGV [24], with HQGA [40] and IGV [24] representing the state of the art. It’s worth noting that we only evaluate methods without large-scale video-text pretraining, excluding pretrained models like MERLOT [48], VIOLET [7], and All-in-one [37]. The exclusion is based on the following reasons: 1) pretrained models require large-scale additional data, making comparisons unfair for methods without pretraining; 2) we aim to highlight the characteristics of our dataset, and the universal knowledge gained by pretrained models could introduce bias in understanding sports scenarios. Table 6 shows the types of visual features used by each model, along with a summary of each method’s insights. For CoMem, HME, and HGA, we use re-implementations provided by [39], and for MASN, HQGA, and IGV, we utilize the official implementations by their respective authors. The baseline model has the same architecture as shown in Fig. 5, with AFT replaced with a standard Transformer Encoder.
In Table 5, the VideoQA methods demonstrate substantial improvements compared to BlindQA. Notably, the most significant improvements are observed for basketball (approximately 40% increase in accuracy), indicating the critical role of visual information in basketball for question answering. Conversely, the improvements for football are less pronounced (about 4%), suggesting that BlindQA already achieves considerable accuracy in this sport. The results suggest high correlations between questions and answers in football, posing challenges for VideoQA models to fully exploit visual information, potentially due to the complexity introduced by a larger number of players.
Examining different question types, the greatest improvements are observed for descriptive questions (around 27% increase in accuracy). Additionally, VideoQA performance on causality and counterfactual questions shows significant improvements compared to BlindQA. However, results for temporal questions are similar between VideoQA methods and BlindQA (except for our AFT), suggesting that VideoQA models may struggle with modeling fine-grained temporal dependencies among actions in sports videos.
@@ 5.4 Further Analysis
Generalization Ability across Sports. In this study, our primary objective is to assess the generalization capabilities across different sports domains. Our approach involves an initial pretraining phase on a set of four sports, followed by fine-tuning on a fifth, distinct sport. We meticulously evaluate the performance of this pretraining-fine-tuning methodology in comparison to training a model entirely from scratch. The outcomes, detailed in Table 7, showcase a notable enhancement in accuracy when employing the pretraining strategy on diverse sports. These findings suggest that the model acquires a nuanced understanding of common semantic features during the initial pretraining. Moreover, the model demonstrates its ability to effectively generalize this acquired knowledge to previously unseen sports during the subsequent fine-tuning phase.
@@
Table 8: The accuracy (%) of various focus lengths and their combinations. Note that the combination of the same focal length implies the use of only one focus length.
@@
1 58.2 / / 58.9 59.0 58.7 58.5
3 / 58.7 / 59.0 59.1 58.6 58.9
5 / / 58.6 58.8 58.9 58.7 58.8
7 / / / 58.7 / / 58.9
9 / / / / 58.8 / 59.0
11 / / / / / 58.4 58.6
80 / / / / / / 58.0
[…]
Impact of Focus Length. In this experiment, we investigate the impact of different focus lengths and their combinations. For short-term dependency, we consider lengths of 1, 3, and 5. For the mid-term, we explore lengths of 7, 9, and 11. Regarding long-term (global) dependency, we set the focus length to the video length, which is 80 seconds. The results are presented in Table 8. From the results, it is evident that the model with global attention performs less effectively compared to those with local attention. However, upon applying our auto-focus attention mechanism, there is a significant improvement in performance. Notably, the combination of focus lengths 3 and 9 achieves the highest accuracy. In our final model, we incorporate the global focus, resulting in the best accuracy at 59.2%.

### v5

@@ 3.3 Diversity, Debiasing and Problem Setting
Diversity of Questions. The limited diversity in generating questions using pre-designed templates may not significantly impact the significance of datasets for several reasons. First, in the context of sports, there is a high level of restriction imposed by rules and specific content, resulting in a more focused attention from the audience on particular actions and events. This inherent structure in sports content narrows down the range of potential questions, making template-based question generation a feasible approach. Secondly, the practice of template-based question generation is widely accepted and employed in the construction of VideoQA datasets [jang2017tgif, xu2017video, yang2021just]. The adoption of this approach in reputable datasets attests to its effectiveness and acceptance. Therefore, while the questions are template-based, their relevance and significance in the context of sports are preserved, ensuring the robustness of the dataset for analysis and evaluation.
@@
Problem Setting. Following the approach in [jang2017tgif, xu2017video], we define an open-ended task based on the generated QA pairs. It is important to note that the open-ended setting, along with the multiple-choice setting, is widely used in the field of VideoQA. We have chosen to commit to the open-ended task instead of the multiple-choice one for two reasons: 1) The open-ended task is more challenging as it requires models to choose from a large answer set rather than selecting from several given options. 2) Limiting the answer choices to only several options would significantly decrease the diversity of the dataset. Specifically, all answers in the QA pairs form an answer pool, which is treated as 191 classes (after discarding classes with fewer than 30 samples).
Interest of Different Participants. We would like to emphasize that the questions in our dataset are intentionally designed to reflect the information needs of different participants, including audience members, commentators, referees, and coaches. For example, we include questions about the outcomes of basketball shots and football attempts, which are typically of interest to both the audience and commentators. We also design questions about the causes or intentions behind specific actions, which can assist coaches in training and instruction. In addition, we include questions related to counting technical movements in gymnastics, which are relevant to referees for scoring purposes. In summary, the question design is grounded in the real interests of these different participants.
@@ 3.4 Dataset Statistics
Additionally, we detail the statistics regarding the temporal and spatial characteristics of each question in Table 4. Specifically, we analyzed each question along two key dimensions: whether it contains spatial directional terms (e.g., “left”/“right”) and whether it includes temporal relational terms (e.g., “before”/“after”). As illustrated in the table, the distribution of questions is balanced with respect to both spatial and temporal attributes, indicating no inherent bias in these aspects.
Table 4: Statistics regarding the temporal and spatial characteristics in Sports-QA.
Before After Other Sum
Left 893 584 7,382 8,859
Right 146 114 7,832 8,092
Other 19,147 18,809 39,166 77,122
Sum 20,186 19,507 54,380 94,073
@@ 4 Auto-Focus Transformer for Sports Video Question Answering
A defining challenge in sports VideoQA is the need to reason over temporal dependencies of diverse scales involving multiple objects. For instance, a question asking about the number of actions requires the model to capture global dependencies in the video, whereas a question querying the effect of an action relies on short-term temporal information. The scale of temporal dependency required varies based on the question. However, current Transformer-based or GNN-based VideoQA methods tend to focus on global dependencies, irrespective of the required scale of temporal information. To address this limitation, we propose the Auto-Focus Transformer (AFT), a new Transformer encoder featuring a novel multi-head Auto-Focus Attention (AFA) mechanism, designed specifically for sports VideoQA.
@@
Specifically, we start by extracting the appearance feature and motion feature of the frame using pretrained ResNet [he2016deep] and I3D [carreira2017quo] (refer to Section 5.1 for more details). Subsequently, we concatenate these features and input them into a linear layer for dimension reduction and feature fusion. Consequently, for each video, we obtain a sequence of frame representations ($\bm{R}$). Regarding the question, we utilize a bi-directional RNN or Transformer to obtain the text representations, encompassing both the global one ($\bm{w}$) and the local ones ($\bm{W}$).
We then capture the dependencies among frames using AFT where the AFA mechanism is computed as follows (the multi-head AFA is defined likewise): given the key set $\left\{\bm{k}_{i}\right\}$, the value set $\left\{\bm{v}_{i}\right\}$, and a query $\bm{q}_{j}$ ($\bm{q}_{j},\bm{k}_{i},\bm{v}_{i}\in\mathbb{R}^{d}$ are mapped from $\bm{R}$ in AFA, and $d$ is the hidden dimension),
@@
where $\mathcal{F}$ is a set of pre-defined focal lengths. $\mathcal{D}_{j}^{f}=\{i||i-j|\leq f\}$ is the index set of the neighbours of $\bm{q}_{j}$ within focal length $f$. $\alpha_{f}\in(0,1)$ is the dynamic weight of focus $f$, satisfying $\sum_{f}\alpha_{f}=1$ and depended on the question. As shown in Fig. 5, $\bm{\alpha}=\{\alpha_{f}\}_{f=1}^{\mathcal{|F|}}$ are obtained by applying linear projection [^6] We also explored a more complex alternative structure—a two-layer MLP—and found that it yields results comparable to those of the linear projection. and softmax to the global representation of question. Since the required scale of temporal dependency is inherently question-dependent, our AFA is designed to selectively focus on only the relevant scale(s) dictated by the question. This targeted modeling ensures the model prioritizes crucial information and filters out irrelevant temporal cues, thereby enhancing the accuracy of predictions.
We denote the output of AFT as $\bm{M}$.
For the video-text integration, we draw inspiration from [li2022invariant] and [jiang2020reasoning], employing a graph convolutional network (GCN) and block fusion [ben2019block]. Refer to [li2022invariant] for detailed information. Ultimately, we obtain a fused feature $\bm{s}$ and predict the answer using a linear projection.
@@ 5.1 Experimental Settings
Pre-processing and Representation. The videos in our dataset are down-sampled to 5 FPS to reduce temporal redundancy. We extract two types of visual features for experiments: appearance feature and motion feature. For each feature type, we obtain both global and local features. Specifically, we detect 10 objects in every two frames of each video using Faster-RCNN [ren2015faster] (with ResNet-101 [he2016deep] as the backbone and a score threshold set to 0.6), pretrained on COCO [lin2014microsoft]. These detected objects serve as local appearance features in our work. Additionally, we use the global average pooling of Res4 activation in Faster-RCNN as the global appearance feature. To extract motion features, we employ I3D [carreira2017quo] pretrained on Kinetics [carreira2017quo]. Specifically, for each frame, we combine it with its adjacent 7 frames, creating a clip that we then feed to I3D to extract a 3D feature map. The global motion feature of a frame is obtained by applying global average pooling to the 3D feature. The local motion feature of an object in a frame is obtained by applying RoIAlign [he2017mask] (output size $1\times 1$) to the 3D feature with its corresponding bounding box. Note that our method uses only the global feature. For language, we explore two types of word embeddings: GloVe [pennington2014glove] and pre-trained BERT [devlin2018bert].
Implementation Details. We set the attention focus to $\mathcal{F}=\{3,9,80\}$ to cover short-term, mid-term, and long-term dependencies, respectively. The hidden dimension $d$ is set to 512. The loss function is cross-entropy. The models are trained for 50 epochs with a learning rate of $1\times 10^{-4}$ and a batch size of 16, using the Adam optimizer [kingma2014adam].
Evaluation Metrics. For evaluation metrics, since we structure the task as a multi-class classification problem, we employ the accuracy of multi-class classification to showcase the models’ performance on our dataset. Additionally, recognizing the long-tail distribution of our dataset, we evaluate methods using F1-score. F1-score is particularly important for sports VideoQA as it explicitly reflects performance averaged over imbalanced classes.
Table 5: Baseline results on Sports-QA, including random choosing, semantic-aware (S-A) random choosing, semantic-aware (S-A) most-frequent-answer choosing, LSTM-based BlindQA, and Transformer-based BlindQA. The best and the second-best results (%) are bold and underlined.
@@
Random S-A Random S-A Frequent LSTM Trans. LSTM Trans.
Accuracy 0.5 24.6 26.8 44.0 43.7 43.9 43.9
F1-score 0.3 5.4 5.7 13.5 14.9 13.8 15.8
@@
CoMem [gao2018motion] 72.8 63.1 69.8 70.0 50.3 77.0 32.9 50.0 60.9 57.1 23.1
HME [fan2019heterogeneous] 64.8 63.3 68.5 71.6 49.3 76.5 31.2 50.4 54.3 56.0 23.6
HGA [jiang2020reasoning] 71.3 65.4 72.2 72.2 51.3 78.1 34.4 52.0 59.2 58.3 24.5
HQGA [xiao2022video] 72.8 64.6 70.5 71.6 51.0 77.6 33.8 54.0 58.2 57.9 25.0
IGV [li2022invariant] 68.8 66.0 70.6 72.4 52.1 78.2 34.7 52.1 59.2 58.5 24.8
MASN [seo2021attend] 71.2 64.2 70.1 66.9 50.5 76.4 33.6 50.2 58.9 57.0 23.4
@@
CoMem [gao2018motion] 74.1 64.0 68.7 70.7 49.3 77.2 32.1 48.4 55.6 56.6 26.4
HME [fan2019heterogeneous] 64.7 63.4 66.2 70.6 49.3 76.6 30.7 46.6 53.6 55.6 22.9
HGA [jiang2020reasoning] 71.6 66.5 69.6 70.7 51.3 78.0 33.8 54.3 56.6 58.1 25.1
HQGA [xiao2022video] 74.1 63.5 67.2 70.8 50.3 76.6 33.6 49.1 57.9 57.0 25.0
IGV [li2022invariant] 72.0 63.9 70.4 72.4 51.7 78.0 34.2 51.9 63.2 58.2 23.8
MASN [seo2021attend] 73.0 63.0 69.7 66.6 50.1 76.4 32.5 50.6 61.2 56.6 24.1
@@
CoMem [gao2018motion] CVPR’18 Mem
HME [fan2019heterogeneous] CVPR’19 Mem
HGA [jiang2020reasoning] AAAI’20 GNN
IGV [li2022invariant] CVPR’22 IG
HQGA [xiao2022video] AAAI’22 HL
MASN [seo2021attend] ACL’21 GNN
@@ 5.2 Baseline Study
Semantic-Aware Most-Frequent-Answer Choosing: Like the Semantic-Aware (S-A) Random Choosing strategy, this approach operates within the constraint of question type. The key distinction lies in its selection criterion: instead of random sampling, it prioritizes the most frequently occurring answer within each question type.
BlindQA: Models are constructed with a question encoder and an MLP-based answer decoder, without exploiting visual information. Two types of word embeddings (Glove and pre-trained BERT) and two question encoders (LSTM and Transformer [vaswani2017attention] [^7] We also experimented with BERT [devlin2019bert], RoBERTa [liu2019roberta], and DeBERTa [hedeberta], as the encoder for the question encoder. However, the improvement brought by them over the baselines was marginal. Considering the balance between accuracy and model complexity, we decided not to include them in our final framework.) are studied.
The results of the baselines are shown in Table 5. Random choosing exhibits poor accuracy and F1-score, but considering constraints from the questions significantly improves performance. The semantic-aware most-frequent-answer strategy yields modest improvements in both accuracy and F1-score. Importantly, however, these performance metrics remain relatively low. This evidence strongly indicates that answer bias in our dataset is well-controlled, as even strategies explicitly leveraging frequency patterns within question types fail to achieve artificially inflated performance. This finding reinforces the robustness of our dataset design, ensuring that model performance reflects genuine understanding rather than reliance on spurious answer distributions. BlindQA achieves considerable results in both accuracy and F1-score, indicating noticeable semantic correlations between questions and answers. By maximizing the likelihood of the answer conditioned on a specific question type, the model fits the answer distribution given different questions in the training set.
@@ 5.3 Benchmark on Sports-QA
We benchmark several VideoQA methods on our Sports-QA dataset, comparing them to CoMem [gao2018motion], HME [fan2019heterogeneous], HGA [jiang2020reasoning], MASN [seo2021attend], HQGA [xiao2022video], and IGV [li2022invariant], with HQGA [xiao2022video] and IGV [li2022invariant] representing the state of the art. It’s worth noting that we only evaluate methods without large-scale video-text pretraining, excluding pretrained models like MERLOT [zellers2021merlot], VIOLET [fu2021violet], and All-in-one [wang2022all]. The exclusion is based on the following reasons: 1) pretrained models require large-scale additional data, making comparisons unfair for methods without pretraining; 2) we aim to highlight the characteristics of our dataset, and the universal knowledge gained by pretrained models could introduce bias in understanding sports scenarios. Table 7 shows the types of visual features used by each model, along with a summary of each method’s insights. For CoMem, HME, and HGA, we use re-implementations provided by [xiao2021next], and for MASN, HQGA, and IGV, we utilize the official implementations by their respective authors. The baseline model has the same architecture as shown in Fig. 5, with AFT replaced with a standard Transformer Encoder.
In Table 6, the VideoQA methods demonstrate substantial improvements compared to BlindQA. Notably, the most significant improvements are observed for basketball (approximately 40% increase in accuracy), indicating the critical role of visual information in basketball for question answering. Conversely, the improvements for football are less pronounced (about 4%), suggesting that BlindQA already achieves considerable accuracy in this sport. The results suggest high correlations between questions and answers in football, posing challenges for VideoQA models to fully exploit visual information, potentially due to the complexity introduced by more players.
Examining different question types, the greatest improvements are observed for descriptive questions (around 27% increase in accuracy). Additionally, VideoQA performance on causality and counterfactual questions shows significant improvements compared to BlindQA. However, results for temporal questions are similar between VideoQA methods and BlindQA (except for our AFT), suggesting that VideoQA models may struggle with modeling fine-grained temporal dependencies among actions.
@@ 5.4 Further Analysis
Generalization Ability across Sports. In this study, our primary objective is to assess the generalization capabilities across different sports domains. Our approach involves an initial pretraining phase on a set of four sports, followed by fine-tuning on a fifth, distinct sport. We meticulously evaluate the performance of this pretraining-fine-tuning methodology in comparison to training a model entirely from scratch. The outcomes, detailed in Table 8, showcase a notable enhancement in accuracy when employing the pretraining strategy on diverse sports. These findings suggest that the model acquires a nuanced understanding of common semantic features during the initial pretraining. Moreover, the model demonstrates its ability to effectively generalize this acquired knowledge to previously unseen sports during the fine-tuning phase.
@@
Table 9: The accuracy (%) of various focal lengths and their combinations on the validation set. Note that the combination of the same focal length implies the use of only one focal length.
@@
1 58.0 / / 58.8 59.1 58.8 58.3
3 / 58.6 / 59.2 59.3 58.3 58.6
5 / / 58.3 59.0 58.7 58.8 58.9
7 / / / 58.4 / / 58.8
9 / / / / 59.0 / 59.2
11 / / / / / 58.2 58.8
80 / / / / / / 58.2
Impact of Focal Length. In this experiment, we investigate the impact of different focal lengths and their combinations on the validation set. For short-term dependency, we consider lengths of 1, 3, and 5. For the mid-term, we explore lengths of 7, 9, and 11. Regarding long-term (global) dependency, we set the focal length to the video length, which is 80 seconds. The results are presented in Table 9. From the results, it is evident that the model with global attention performs less effectively compared to those with local attention. However, upon applying our auto-focus attention mechanism, there is a significant improvement in performance. Notably, the combination of focal lengths 3 and 9 achieves the highest accuracy. In our final model, we incorporate the global focus, resulting in the best accuracy at 59.2%.
Performance of Multi-modal Large Language Models (MLLMs). Following the setup outlined in [yang2024sports], we evaluated the zero-shot performance of various multimodal large language models (MLLMs) on our dataset. The corresponding results from [yang2024sports] are presented in Table 10. As shown in the table, the lower scores reveal that MLLMs still have substantial scope for advancement in video-based sports recognition. This outcome underscores that video-based sports understanding remains a notable challenge for MLLMs, emphasizing the intricate nature of multimodal sports comprehension tasks. Since video-based sports recognition is primarily rooted in identifying specific sports and their associated actions, MLLMs require targeted enhancements in this domain to improve performance. We also observe that the CoT (Chain-of-Thought) variants of MiniGPT and ChatUniVi perform worse than their original baselines. This degradation likely arises because these models were not pretrained to generate explicit reasoning steps, and introducing CoT changes the expected output distribution from concise answers to verbose reasoning chains. The additional reasoning text often introduces noise or hallucinated visual details, which can propagate errors to the final prediction. Moreover, since our benchmark assess only short final answers, the longer CoT outputs may be penalized. To further assess the difficulty of Sports-QA, we fine-tuned an advanced multi-modal LLM, Qwen2.5-VL-3B [bai2025qwen2], on our dataset. As shown, the zero-shot model performs poorly due to the large domain gap, while fine-tuning leads to a substantial accuracy improvement, reaching 63.71%. Despite this improvement, the fine-tuned model still leaves considerable room for further progress, highlighting the challenging nature of our benchmark.
Table 10: The accuracy (%) of different MLLMs (zero-shot or finetuning setting) on Sports-QA.
Model Desc. Temp. Causal Counter. Acc.
Minigpt-4-7B [zhu2023minigpt] 36.85 14.10 23.20 36.51 26.70
Minigpt-4-7B (CoT) 36.38 14.15 19.79 34.54 26.26
Chat-UniVi-7B [jin2024chat] 45.31 16.23 24.23 6.58 31.52
Chat-UniVi-7B (CoT) 45.04 14.71 23.09 13.82 30.81
Video-LLaVA-7B [zhu2023languagebind] 42.80 13.52 36.80 42.76 30.33
Video-LLaVA-7B (CoT) 43.46 13.50 35.57 39.47 30.55
LLaVA-Video-Qwen2-7B [zhang2024video] 45.77 15.94 37.73 41.78 32.73
Qwen2.5-7B [bai2025qwen2] 46.57 16.43 38.41 42.87 33.40
Qwen2.5-3B 37.81 15.45 25.61 37.57 27.77
Qwen2.5-3B-FT 81.71 42.63 58.11 58.97 63.71